Find the probability of drawing 3 queens in a row:
a. if the card is replaced after each draw
b. if the card is not replaced after each draw

Answer by stanbon(75887)

Find the probability of drawing 3 queens in a row:
a. if the card is replaced after each draw
Ans: (4/52)^3 = 0.0004552...
----------------------------------------
b. if the card is not replaced after each draw
Ans: (4/52)(3/51)(2/50) = 0.0001810..
------------------
Another way to get that answer:
4C3/52C3 = 0.0001810..
=============================
Cheers,
Stan H.